# FissionVAE: Federated Non-IID Image Generation with Latent Space and Decoder Decomposition The source code of FissionVAE is publicly available at: https://github.com/Rand2AI/FissionVAE

### v1

@@ Abstract
Federated learning is a machine learning paradigm that enables decentralized clients to collaboratively learn a shared model while keeping all the training data local. While considerable research has focused on federated image generation, particularly Generative Adversarial Networks, Variational Autoencoders have received less attention. In this paper, we address the challenges of non-IID (independently and identically distributed) data environments featuring multiple groups of images of different types. Specifically, heterogeneous data distributions can lead to difficulties in maintaining a consistent latent space and can also result in local generators with disparate texture features being blended during aggregation. We introduce a novel approach, FissionVAE, which decomposes the latent space and constructs decoder branches tailored to individual client groups. This method allows for customized learning that aligns with the unique data distributions of each group. Additionally, we investigate the incorporation of hierarchical VAE architectures and demonstrate the use of heterogeneous decoder architectures within our model. We also explore strategies for setting the latent prior distributions to enhance the decomposition process. To evaluate our approach, we assemble two composite datasets: the first combines MNIST and FashionMNIST; the second comprises RGB datasets of cartoon and human faces, wild animals, marine vessels, and remote sensing images of Earth. Our experiments demonstrate that FissionVAE greatly improves generation quality on these datasets compared to baseline federated VAE models.
@@ 1 Introduction
Generative models have attracted increasing attention in recent years due to their impressive ability to generate new data across various modalities, including images, texts, and audios [12, 28, 2]. As these models, like other deep learning systems, require substantial amounts of data, concerns regarding data privacy have elevated among regulatory authorities and the public. Unlike the traditional centralized learning paradigm, which collects all data on a single computer system for training, federated learning allows private data to remain on the owner’s device. In this paradigm, local devices train models independently, and a central server aggregates these models without accessing the individual data directly. Although this distributed approach enhances privacy protection, it also introduces unique challenges not encountered in centralized systems. Since data remains distributed across various client devices, the training samples are not guaranteed to be identically distributed. This can lead to inconsistencies in learning objectives among clients, resulting in degraded performance when these models are aggregated on the server.
@@
In the context of FL with non-IID data, generative models such as Generative Adversarial Networks (GANs) [6] and Variational Autoencoders (VAEs) [16] face additional challenges. These models involve sampling from a latent distribution, and the generator or decoder trained on client devices may develop differing interpretations of the same latent space. This discrepancy can lead to difficulties in maintaining a consistent and unified latent space, resulting in overlapping representations across models, shown in Figure 2. A further challenge arises from the role of the generator or decoder, which are tasked with mapping latent inputs to the sample space by synthesizing the shape, texture, and colors of images. Aggregating generative models trained on non-IID image data can produce artifacts that appear as a blend of disparate image types, because generators trained on non-IID local data capture the characteristics of varied visual features (shown in Figure 4). Specifically for GANs, another problem arises from local discriminators, which may provide conflicting feedback that hinders model convergence. With the limited data available in FL settings, discriminators can quickly overfit to the training samples [14]. If an updated generator from the server produces images of classes not present in a client’s local dataset, the local discriminator might incorrectly label well-generated images as fake, simply because they do not match the local data distribution. This mislabeling can significantly impede the generator’s ability to synthesize realistic images.
@@
Existing research on generative models for non-IID data in federated learning (FL) has primarily focused on GANs. MDGAN [8] proposes exchanging local discriminators among clients during training. This strategy allows discriminators to access a broader spectrum of local data, thereby avoiding biased feedback to the generator. The authors of [34] uses the local discriminator that gives the highest score to a generated sample to update the global generator, promoting the idea that local discriminators should only judge samples from familiar distributions. In [33], the authors aggregate generators at the group level for client groups sharing similar data distributions before performing a global aggregation, then the global generator is aggregated similar to [34]. Both [34] and [33] involve sending synthesized samples back to local clients, which could potentially increase the risk of compromising client data privacy.
@@
Studies employing VAEs solely for image generation purposes are less common. The works in [3] and [9] utilize VAEs to produce synthetic images that assist in training global classifiers. In [3], the global decoder generates minority samples for local classifiers by sampling from class means with added noise. The approach in [9] treats converged local decoders as teacher models and uses knowledge distillation to train a global generator on the server side without further local updates. While this decoder can produce useful samples for classification tasks, it risks overfitting to the potentially flawed output from local decoders and lacks generative diversity, which is crucial for high-quality image generation. Recent studies [1, 25] have shown that generative models trained on generated samples instead of real data are prone to collapsing.
@@
In response to the challenges posed by non-IID data in federated image generation, we introduce a model named FissionVAE. This model is specifically tailored to environments featuring multiple groups of images of different types. To mitigate the problem of mixed latent space interpretation, FissionVAE decomposes the latent space into distinctive priors, effectively managing the diversity of data distributions across different image types. We further refine this approach by investigating strategies for encoding the prior Gaussians. Additionally, to prevent the blending of unrelated visual features in the generated outputs, FissionVAE employs tailored decoder branches for each client group. This method not only accommodates the unique characteristics of each data subset but also enhances the model’s generative capabilities in highly heterogeneous environments. The primary contributions of our research are detailed as follows:
@@
The remainder of the paper is organized as follows: Section II reviews existing work related to generative models in federated learning. In Section III, we begin by outlining the preliminaries of Federated Averaging, followed by a detailed discussion of the baseline FedVAE model and the FissionVAE variants we propose. Section IV presents the experimental setup, including the configuration details and an analysis of the results. Finally, we conclude the paper in Section V with a summary of our findings and a discussion on potential future directions.
2 Preliminaries
2.0.1 Federated Averaging
Federated Averaging (FedAvg) [19] is the most commonly used algorithm for federated learning tasks. In FedAvg, the central server first distribute the initialized model to clients, ensuring that all clients starts local training with the identical model parameters. Clients train their models with local data for specified local epochs, then return their models to the central server. The central server aggregates collected client models by taking the weighted average of these models:
$\omega_{r}=\sum_{i=1}^{|S|}{\frac{{n_{i}}}{N_{S}}\omega_{i}}$ (1)
where $|S|$ is the number of clients that participate current training round, $n_{i}$ is the number of training samples on client $i$ while $\omega_{i}$ is the model parameters for this client, $\omega_{r}$ is the parameters of the global model, and $N_{S}$ is the total number of training samples across clients in current round. Once aggregated, the central server distribute the updated model to clients in the next training round. This process is repeated until the global model converges.
2.0.2 Variational Autoencoders
VAEs are a class of generative models designed to approximate the probability $p(x)$ of the observed data $x$, which is typically intractable. Central to the optimization of VAEs is the concept of evidence lower bound (ELBO), which provides a tractable objective by bounding the logarithm of $p(x)$ through Jensen’s inequality. The ELBO is defined as follows:
[…]
where $q_{\phi}(\cdot)$ is the encoder parameterized by $\phi$, and $z$ denotes the latent code derived from the input data $x$. The ELBO can be further decomposed into the following terms by splitting the logarithm,
$\rm ELBO=\mathbb{E}_{q_{\phi}(z|x)}[\log p_{\theta}(x|z)]-D_{KL}(q_{\phi}(z|x)||p(z))$ (3)
Here, $p_{\theta}(\cdot)$ is the decoder parameterized by $\theta$. The first term of the ELBO represents the expected log-likelihood of the data reconstructed from the latent code $z$. The second term is the Kullback-Leibler divergence between the encoded distribution $q_{\phi}(z|x)$ and a predefined prior distribution $p(z)$.
During training, the encoder $q_{\phi}$ estimates the latent code $z$ based on the input data $x$, aiming to align this estimate with the prior distribution $p(z)$. The decoder $p_{\theta}$ then attempts reconstruct the original input $x$ from the estimated $z$. The decomposition above indicates that training objective for VAEs is optimizing both the quality of data reconstruction and adherence to the latent space structure. In the generation phase, new data instances are produced by sampling from the prior distribution $p(z)$, which are then transformed solely by the decoder into the data space.
2.0.3 Variational Autoencoders in Federated Learning
Following the discussion on the ELBO, we note that in a centralized setting, we maximize the expectation of the reconstruction likelihood across all data, as presented in Equation 3. In contrast, federated learning distributes the training samples $x$ are split into subsets $\{x_{1}\cup x_{2}\cup...\cup x_{n}\}$ across $n$ clients. A common assumption in federated learning is that client data are not identically distributed, yet they are independent. Furthermore, we assume that client data are conditionally independent given the latent codes $z$, such that $p(x_{i}|z,x_{j})=p(x_{i}|z),i\neq j$. This leads to the following factorization,
$p(x_{1},...,x_{n}|z)$ $=p(x_{1}|z)p(x_{2}|x_{1},z)p(x_{3},...,x_{n}|z)$ (4)
$=\Pi_{i=1}^{n}p(x_{i}|z)$
Therefore, the reconstruction likelihood in the ELBO under the federated assumption can be expressed as:
$\log p_{\theta}(x|z)=\Sigma_{i=1}^{n}\log p_{\theta}(x_{i}|z)$ (5)
Equation 5 suggests that when training VAEs within the federated learning framework, we can independently maximize the ELBO on each client. Local VAEs are then aggregated through algorithms such as FedAvg to maximize the global ELBO.
Regarding existing research on federated VAEs for image generation, the work in [5] shows that VAEs can generate handwritten digit images within an FL framework and [20] focuses on incorporating differential privacy into the training process of federated VAEs for image generation. In [3], the authors integrate a federated VAE into the main classification task to help clients balance their local data. This involves training a global VAE where the encoder finds the average latent representation of local classes, these representations are then sent to the central server. When needed, the global decoder and collected latent representations are distirbuted to clients for image generation. The work in [9] aims at one-shot federated learning for classification tasks, where the server distributes the initialized model and only collects the converged local models without further updates. Here, the VAE is used as a data generator to create a balanced training set through knowledge distillation from local decoders. As discussed in Section I, while effective for classification tasks where balancing the training dataset is paramount, the lack of real data exposure and a cohesive latent space constructed by an encoder challenges the utility of this approach in producing high-quality images.
Table 1: Summary of Explored Strategies for Non-IID Image Generation with VAE
Approach
Latent Space
Decomposition
Decoder Branch
Decomposition
Group-specific
Aggregation
Hierarchical
Inference
FedVAE (baseline)
FissionVAE + L
FissionVAE + L + D
FissionVAE + L + D + H
@@ 3 Investigating Strategies for Non-IID Image Generation with VAEs
In this section, we describe our methodology for exploring VAE configurations tailored for generating images under non-IID conditions in a federated learning framework. We specifically address scenarios where clients are categorized based on distinct data distributions. For illustrative purposes, we consider a case where some clients exclusively possess hand-written digits from the MNIST dataset, while others maintain only clothing images from the FashionMNIST dataset. We follow to the standard federated learning framework, wherein a central server is tasked with aggregating updates from the clients and subsequently distributing the updated model back to them. Each client retains a subset of data representative of its respective group and conducts local training independently. A more complex scenario with RGB images and a larger number of client groups is explored and discussed in the experimental section (Section IV). We summarize the explored strategies in Table 1.
@@ 3.1 FedVAE
A straightforward strategy for implementing VAEs in federated learning is using a unified encoder-decoder architecture. In this configuration, all clients share a common latent space (often predefined as the normal distribution $\mathcal{N}(0,1)$) and the central server indiscriminately aggregates client models at the end of each training round. Fig. 1 illustrates this training scheme.
@@
Figure 2: Qualitative Results of baseline FedVAE on the Mixed MNIST dataset. Both reconstructed and generated output shows the artifact that blends the visual features from distinct datasets. The t-SNE visualization shows that the latent representations of test samples are also overlapped.
Despite the simplicity of this strategy, it present significant challenges in the non-IID scenario. Specifically, employing a single prior distribution for the latent space does not account for the distinct data distributions across different clients. Encoders from different client groups may map their uniquely distributed data into the same region of the latent space. Consequently, client decoders might interpret this shared latent space differently, leading to inconsistencies or even conflicts among client models during aggregation at the server. Figure 2 shows randomly selected reconstruction and generation samples produced after training the federated Vanilla VAE on the combined dataset of MNIST and FashionMNIST. These samples clearly exhibit artifacts that appear to blend features of handwritten digits with clothing items, indicating the aggregation conflicts inherent in this method.
@@ 3.2 FissionVAE with Latent Space Decoupling
To address the conflicting latent space issue identified above, we propose decomposing the latent space according to different data groups, while maintaining a unified architecture for the encoder and decoder. This approach corresponds to FissionVAE+L in Table 1 and its architecture is shown in Fig. 3.
Figure 3: An illustration of FissionVAE with Latent Space Decompostion. The latent variables are required to follow their respective group prior distributions. The model is aggregated the same way as the baseline FedVAE.
Specifically, the encoder maps the input data to different distributions based on the client’s group. For instance, MNIST client may map to $\mathcal{N}(-1,1)$ and FashionMNIST clients to $\mathcal{N}(-1,1)$. The KL divergence in the ELBO for this model is given by:
@@
Figure 4: Qualitative results of federated FissionVAE with latent space decomposition on the Mixed MNIST dataset. Despite improved reconstruction quality and separated latent space, feature blending artifacts can be still observed in generated samples.
Figure 4 shows reconstruction and generation samples produced after training the FissionVAE with latent space decomposition on the Mixed MNIST dataset. While the quality of reconstructed images are greatly improved compared to the baseline FedVAE, the generated images still exhibit a mixture of handwritten digits and clothing items, even when explicitly sampling from their respective latent distributions. This suggests that while decomposing latent encoding helps improving reconstructions, the unified decoder still blends features due to the aggregation of model weights from diverse visual domains. This observation motivates the architecture described in the next section, where the decoder is also split based on client groups.
@@ 3.3 FissionVAE with Group-specific Decoder Branches
Non-Hierarchical FissionVAE Building on the concept introduced by FissionVAE with latent space decomposition, we further refines non-IID data generation by incorporating decoder branches specific to each data group while maintaining a unified encoder. This design allows the central server to aggregate the encoder updates agnostically of the client groups, whereas decoder branches are aggregated specifically according to their corresponding groups. In addition, this approach also offers flexibility in the choice of the prior latent distribution $p(z)$, which can be identical across groups or defined uniquely for each group to exert more explicit control over the data generation through the decoder. In scenarios where $p(z)$ are identical for all groups, although the encoder projects different samples into the same latent space, group-specific decoders ensure that the latent space is interpreted independently. Nevertheless, we find that the uniquely defined priors result in better generation quality as the number of client groups increases. Figure 5 illustrates this branching architecture which corresponds to FissionVAE+L+D in Table 1.
Figure 5: An illustration of FissionVAE with Decoder Branch Decomposition. This FissionVAE creates decoders specific to client groups and enforces constraints for latent variable priors. The encoder is aggregated across groups while the group-specific decoder is only aggregated from local models within the corresponding group.
Figure 6: Qualitative results of federated FissionVAE with decomposed decoder branches on the Mixed MNIST dataset. Images of different types are more discernable compared to methods above.
Figure 6 shows randomly selected reconstruction and generation samples produced after training the FissionVAE with decoder branches on the combined dataset of MNIST and FashionMNIST. The results indicate a significant reduction in the blending feature issue in previously discussed VAE architectures.
Hierachical FissionVAE Next, we show that the branching architecture can be enhanced by integrating hierarchical inference [15, 27] to the federated learning framework, which enables the use of deeper network structures to capture more complex data distributions. Fig 7 depicts the FissionVAE with two levels of hierarchical inference, which is the FissionVAE+L+D+H configuration in Table 1. In this architecture, the first encoder module estimates $q(z_{1}|x)$ from the input data, then the second encoder module estimates $q(z_{2}|z_{1})$ based on the first level latent code. The decoder reverses the encoding process, which estimates $p(z_{1}|z_{2})$ based on $z_{2}$ to reconstruct $z_{1}$, and subsequently reconstructs the original input $x$ by estimating $p(x|z_{1})$.
Following the convention in hierarchical VAEs, we assume conditional independence among the latent codes, where the encoding and decoding processes can be defined as follows:
$q(z_{1},z_{2}|x)=q_{\phi}(z_{1}|x)q_{\phi}(z_{2}|z_{1})$ (7)
$p(x,z_{1},z_{2})=p(z_{2})p_{\theta}(x|z_{1})p_{\theta}(z_{1}|z_{2})$ (8)
The ELBO for this hierarchical VAE is expressed as,
@@
In the equation above, the first term is the reconstruction term as it is the expectation of the log-likelihood for the input samples under the distribution estimated from the encoded $z_{1}$, the second term is the prior matching term which is enforcing the encoded $z_{2}$ to conform the prior distribution $z_{2}\sim\mathcal{N}(0,1)$, and the last term is the consistency term which requires $z_{1}$ from either the encoder or the decoder to be consistence. In practice, we find that adding the reconstruction loss from $z_{2}$ to $x$ is also crucial for generating meaningful samples. Optionally, perceptual losses such as the VGG loss [18] or the structural similarity index measure (SSIM) [30] loss can be used to promote the fidelity of reconstructed images. However, no significant improvement is observed in our experiments. Therefore no perceptual loss is included in our implementation. The final loss function for the hierarchical and branching FissionVAE then becomes,
@@
where we minimize the KL divergence for $z_{1}$ only when the prior distribution for $z_{1}$ is defined, otherwise this term is zero and the estimation for $z_{1}$ is constrained by the latent matching term in Equation 6.
4 Experiments
4.1 Datasets and Evaluation Metrics
The experiment setup for federated image generation is summarized in Table 2. We constructed two composite datasets to evaluate the performance of the proposed federated VAEs. The first dataset Mixed Mnist combines MNIST [17] and FashionMNIST [32]. These two datasets consists of images of handwritten digits from 0 to 9 and 10 classes of various clothing items, respectively. Images from these two datasets share a common size of $28\times 28$ pixels, and we retain this size for generation and evaluation purposes. In addition, each dataset contains 60,000 training samples and 10,000 test samples. We divided training samples into two client groups, one for each dataset, with each group comprising 10 clients. The training samples were randomly and evenly distributed among the clients within each group. For evaluation, we used the default test samples from both MNIST and FashionMNIST to approximate the true data distribution. We generated an equal number of samples from the global model on the central server to evaluate performance.
In contrast to the Mixed MNIST dataset, which primarily serves proof-of-concept purposes, we constructed a second, more complex and realistic composite dataset named CHARM, encompassing five distinct datasets: Cartoon faces from the Anime Face Dataset [4], Human faces from CelebA-HQ [13], Animals from Animals with Attributes 2 (AwA2) [31] with 50 classes, Remote sensing images of 10 land types from EuroSAT-RGB [10], and Marine vessels from MARVEL [7] with 26 classes. For AwA2 and MARVEL, we used preprocessed square images from Meta-Album [29]. All images in this composite dataset are resized to 32$\times$32 pixels for generation and evaluation. We established 20 clients for each of the five groups in CHARM. We utilized 20,000 randomly selected images from these datasets for training, with an additional 5,000 images from each used for evaluation. As with Mixed MNIST, we generated a number of images equal to the evaluation set using the global model for performance evaluation.
On the Mixed MNIST dataset, the encoder and decoder for all VAEs only consists of Multi-layer Perceptrons (MLP). For models trained on CHARM, the encoder $q(z_{1}|x)$ and the decoder $p(x|z1)$ are primarily convolutional with MLP used for latent variable estimation. The encoder $q(z_{2}|z_{1})$ and the decoder $p(z_{1}|z_{2})$ are MLPs. Bernoulli distribution is used to determine whether or not a client participate in current training round. For the Mixed MNIST dataset, we set the distribution to $B(0.5)$, indicating a 50% chance of participation. For CHARM, we adjusted the participation rate to $B(0.25)$, meaning approximately 25 out of 100 clients participate in each training round. Regarding the hyperparameters for training, we use a learning rate of $1\times 10^{-3}$ for the Mixed MNIST and $1\times 10^{-4}$ for the CHARM dataset. The total number of training rounds is 70 on Mixed MNIST and 500 on CHARM. For both datasets, client update local models for 5 epochs with a batch size of 32. In the centralized setting, the number of training epochs on Mixed MNIST is 70 and 250 for CHARM. Other hyperparamters are kept the same as the federated configuration.
We employed the Fréchet Inception Distance (FID) [11] and Inception Score (IS) [24] to measure the quality of generated images. The negative log likelihood (NLL) of the ELBO was used to reflect the model’s performance on reconstructing test images. We employ Equation 3 to compute NLL for VAEs with no hierarchical structure and use Equation 9 for hierarchical VAEs. The Inception model [26] pre-trained on ImageNet [23] is used to evaluate the IS.
Table 2: Federated Learning Experiment Setup
Dataset Client Groups # Clients Participation Rate Model Type # Local Samples Per Client # Evaluation Samples
Mixed MNIST MNIST[17], FashionMNIST[32] 20 0.5 MLP 6,000 20,000
CHARM Anime Face [4] CelebA [13], AwA2 [31] EuroSet [10], MARVEL [7] 100 0.25 Convolution + MLP 1,000 25,000
Table 3: Evaluation of proposed FissionVAEs on the Mixed MNIST and CHARM dataset. +L is for decoupled latent space. +D is for branching decoders. +H is for the hierarchical architecture.
@@ 4.1 Datasets and Evaluation Metrics
Best results in are in bold. Second best results are underlined. $\uparrow$ denotes the higher the better, while $\downarrow$ means the lower the better.

### v2

@@ Abstract
Federated learning is a machine learning paradigm that enables decentralized clients to collaboratively learn a shared model while keeping all the training data local. While considerable research has focused on federated image generation, particularly Generative Adversarial Networks, Variational Autoencoders have received less attention. In this paper, we address the challenges of non-IID (independently and identically distributed) data environments featuring multiple groups of images of different types. Non-IID data distributions can lead to difficulties in maintaining a consistent latent space and can also result in local generators with disparate texture features being blended during aggregation. We thereby introduce FissionVAE that decouples the latent space and constructs decoder branches tailored to individual client groups. This method allows for customized learning that aligns with the unique data distributions of each group. Additionally, we incorporate hierarchical VAEs and demonstrate the use of heterogeneous decoder architectures within FissionVAE. We also explore strategies for setting the latent prior distributions to enhance the decoupling process. To evaluate our approach, we assemble two composite datasets: the first combines MNIST and FashionMNIST; the second comprises RGB datasets of cartoon and human faces, wild animals, marine vessels, and remote sensing images. Our experiments demonstrate that FissionVAE greatly improves generation quality on these datasets compared to baseline federated VAE models.
@@ 1 Introduction
Generative models have attracted increasing attention in recent years due to their impressive ability to generate new data across various modalities, including images Ho et al. (2020), texts Touvron et al. (2023), and audios Borsos et al. (2023). As these models, like other deep learning systems, require substantial amounts of data, concerns regarding data privacy have elevated among regulatory authorities and the public. Unlike the traditional centralized learning paradigm, which collects all data on a single computer system for training, federated learning allows private data to remain on the owner’s device. In this paradigm, local devices train models independently, and a central server aggregates these models without accessing the individual data directly. Although this distributed approach enhances privacy protection, it also introduces unique challenges not encountered in centralized systems. Since data remains distributed across various client devices, the training samples are not guaranteed to be identically distributed. This can lead to inconsistencies in learning objectives among clients, resulting in degraded performance when these models are aggregated on the server.
@@
In the context of FL with non-IID data, generative models such as Generative Adversarial Networks (GANs) Goodfellow et al. (2014) and Variational Autoencoders (VAEs) Kingma and Welling (2014) face additional challenges. These models involve sampling from a latent distribution, and the generator or decoder trained on client devices may develop differing interpretations of the same latent space. This discrepancy can lead to difficulties in maintaining a consistent and unified latent space, resulting in ambiguous latent representations. A further challenge arises from the role of the generator or decoder, which are tasked with mapping latent inputs to the sample space by synthesizing the shape, texture, and colors of images. Aggregating generative models trained on non-IID image data can produce artifacts that appear as a blend of disparate image types, because generators trained on non-IID local data capture the characteristics of varied visual features. Specifically for GANs, another problem arises from local discriminators, which may provide conflicting feedback that hinders model convergence. With the limited data available in FL settings, discriminators can quickly overfit to the training samples Karras et al. (2020). If an updated generator from the server produces images of classes not present in a client’s local dataset, the local discriminator might incorrectly label well-generated images as fake, simply because they do not match the local data distribution. This mislabeling can significantly impede the generator’s ability to synthesize realistic images.
@@
Existing research on generative models for non-IID data in federated learning (FL) has primarily focused on GANs. MDGAN Hardy et al. (2019) proposes exchanging local discriminators among clients during training. This strategy allows discriminators to access a broader spectrum of local data, thereby avoiding biased feedback to the generator. The authors of Yonetani et al. (2019) uses the local discriminator that gives the highest score to a generated sample to update the global generator, promoting the idea that local discriminators should only judge samples from familiar distributions. In Xiong et al. (2023), the authors aggregate generators at the group level for client groups sharing similar data distributions before performing a global aggregation, then the global generator is aggregated similar to Yonetani et al. (2019). Both Yonetani et al. (2019) and Xiong et al. (2023) involve sending synthesized samples back to local clients, which could potentially increase the risk of compromising client data privacy.
@@
Studies employing VAEs solely for image generation purposes are less common. The works in Chen and Vikalo (2023) and Heinbaugh et al. (2023) utilize VAEs to produce synthetic images that assist in training global classifiers. In Chen and Vikalo (2023), the global decoder generates minority samples for local classifiers by sampling from class means with added noise. The approach in Heinbaugh et al. (2023) treats converged local decoders as teacher models and uses knowledge distillation to train a global generator on the server side without further local updates. While this decoder can produce useful samples for classification tasks, it risks overfitting to the potentially flawed output from local decoders and lacks generative diversity, which is crucial for high-quality image generation. Recent studies Bohacek and Farid (2023) Shumailov et al. (2024) have shown that generative models trained on generated samples instead of real data are prone to collapsing.
@@
In response to the challenges posed by non-IID data in federated image generation, we introduce a model named FissionVAE. This model is specifically tailored to environments featuring multiple groups of images of different types. To mitigate the problem of mixed latent space interpretation, FissionVAE decomposes the latent space into distinctive priors, hence adapting to the diverse data distributions across different image types. We further refine this approach by investigating strategies for encoding the prior Gaussians. Additionally, to prevent the blending of unrelated visual features in the generated outputs, FissionVAE employs specialized decoder branches for each client group. This method not only accommodates the unique characteristics of each data subset but also enhances the model’s generative capabilities in highly heterogeneous environments. The primary contributions of our research are detailed as follows:
@@
The remainder of the paper is organized as follows: In Section 2, we describe the baseline FedVAE model and the FissionVAE variants we propose. Section 3 presents the experimental setup, including the configuration details and an analysis of the results. Finally, we conclude the paper in Section 4 with a summary of our findings and a discussion on potential future directions.
Figure 1: Qualitative results of the baseline FedVAE and proposed FissionVAEs. As we further decoupling the latent space and decoders in the federated environment, the quality of generated images is improved.
@@ 2 Investigating Strategies for Non-IID Image Generation with VAEs
In this section, we describe our methodology for exploring VAE configurations tailored for generating images under non-IID conditions in a federated learning framework. For background on FL and VAEs, please refer to the supplementary material. We specifically address scenarios where clients are categorized based on distinct data distributions. For illustrative purposes, we consider the case where some clients exclusively possess hand-written digit images from the MNIST dataset, while others maintain only clothing images from the FashionMNIST dataset. We follow to the standard federated learning framework, wherein a central server is tasked with aggregating updates from the clients and subsequently distributing the updated model back to them. FedAvg McMahan et al. (2023) is employed for server-side aggregation. Each client retains a subset of data representative of its respective group and conducts local training independently. A more practical scenario with RGB images and a larger number of client groups is explored and discussed in the experiments section (Section 3).
@@ 2.1 FedVAE
A straightforward strategy for implementing VAEs in federated learning is using a unified encoder-decoder architecture. In this configuration, all clients share a common latent space (often predefined as the normal distribution $\mathcal{N}(0,1)$) and the central server indiscriminately aggregates client models at the end of each training round. Fig. 2 illustrates this baseline training scheme.
@@
Despite the simplicity of this strategy, it present significant challenges in the non-IID scenario. Specifically, employing a single prior distribution for the latent space does not account for the distinct data distributions across different clients. Encoders from different client groups may map their uniquely distributed data into the same region of the latent space. Consequently, client decoders might interpret this shared latent space differently, leading to inconsistencies or even conflicts among client models during aggregation at the server. Figure 1 shows randomly generated samples produced after training the federated Vanilla VAE on the combined dataset of MNIST and FashionMNIST. These samples clearly exhibit artifacts that appear to blend features of handwritten digits with clothing items, indicating the aggregation conflicts inherent in this method.
@@ 2.2 FissionVAE with Latent Space Decoupling
To address the conflicting latent space issue identified above, we propose decomposing the latent space according to different data groups, while maintaining a unified architecture for the encoder and decoder. This approach corresponds to the architecture shown in Fig. 3.
Figure 3: An illustration of FissionVAE with Latent Space Decoupling. The latent variables are forced to follow their respective group prior distributions. The model is aggregated the same way as the baseline FedVAE.
When decoupling the latent space, the encoder maps the input data to different distributions based on the client’s group. For instance, MNIST client may map to $\mathcal{N}(-1,1)$ and FashionMNIST clients to $\mathcal{N}(-1,1)$. The KL divergence in the ELBO for this model is given by:
@@
Figure 1 shows randomly generated amples produced after training the FissionVAE with latent space decoupling on the Mixed MNIST dataset. While the quality of reconstructed images are improved compared to the baseline FedVAE, the generated images still exhibit a mixture of handwritten digits and clothing items, even when explicitly sampling from their respective latent distributions. This suggests that while decomposing latent encoding helps improving reconstructions, the unified decoder still blends features due to the aggregation of model weights from diverse visual domains. This observation motivates the architecture described in the next section, where the decoder is also split based on client groups.
@@ 2.3 FissionVAE with Group-specific Decoder Branches
Non-Hierarchical FissionVAE Building on the concept introduced by FissionVAE with latent space decoupling, we further refines non-IID data generation by incorporating decoder branches specific to each data group while maintaining a unified encoder. This design allows the central server to aggregate the encoder updates agnostically of the client groups, whereas decoder branches are aggregated specifically according to their corresponding groups. In addition, this approach also offers flexibility in the choice of the prior latent distribution $p(z)$ for each group to exert more explicit control over the data generation through the decoder. Figure 5 illustrates this branching architecture.
Figure 5: An illustration of FissionVAE with Decoder Branch Decoupling. This FissionVAE creates decoders specific to client groups and enforces constraints for latent variable priors. The encoder is aggregated across groups while the group-specific decoder is only aggregated from local models within the corresponding group.
Figure 1 also includes randomly generated samples produced after training the FissionVAE with decoder branches. The results indicate a significant reduction in the blending feature issue in previously discussed VAE architectures.
Hierachical FissionVAE Next, we show that the branching architecture can be enhanced by integrating hierarchical inference Kingma et al. (2016) Sønderby et al. (2016) to the federated learning framework, which enables the use of deeper network structures to capture more complex data distributions. Fig 4 depicts the FissionVAE with two levels of hierarchical inference. In this architecture, the first encoder module estimates $q(z_{1}|x)$ from the input data, then the second encoder module estimates $q(z_{2}|z_{1})$ based on the first level latent code. The decoder reverses the encoding process, which estimates $p(z_{1}|z_{2})$ based on $z_{2}$ to reconstruct $z_{1}$, and subsequently reconstructs the original input $x$ by estimating $p(x|z_{1})$.
Following the convention in hierarchical VAEs, we assume conditional independence among the latent codes.Then the ELBO for this hierarchical VAE is expressed as (refer to supplementary material for derivation),
@@
In the equation above, the first term is the reconstruction term as it is the expectation of the log-likelihood for the input samples under the distribution estimated from the encoded $z_{1}$, the second term is the prior matching term which is enforcing the encoded $z_{2}$ to conform the prior distribution $z_{2}\sim\mathcal{N}(0,1)$, and the last term is the consistency term which requires $z_{1}$ from either the encoder or the decoder to be consistence. In practice, we find that adding the reconstruction loss from $z_{2}$ to $x$ is also crucial for generating meaningful samples. Optionally, perceptual losses such as the VGG loss Ledig et al. (2017) or the structural similarity index measure (SSIM) Wang et al. (2004) loss can be used to promote the fidelity of reconstructed images. However, no significant improvement is observed in our experiments. Therefore no perceptual loss is included in our implementation. The final loss function for the hierarchical and branching FissionVAE then becomes,
@@
Here we minimize the KL divergence for $z_{1}$ only when the prior distribution for $z_{1}$ is explicitly defined, otherwise the model learns the latent distribution by itself.
@@
Table 1: Evaluation of proposed FissionVAEs on the Mixed MNIST and CHARM dataset. +L is for decoupled latent space. +D is for branching decoders. +H is for the hierarchical architecture. Best results in are in bold. Second best results are underlined. $\uparrow$ denotes the higher the better, while $\downarrow$ means the lower the better.
3 Experiments
3.1 Datasets and Evaluation Metrics
We evaluated the proposed federated VAEs using two composite datasets. Mixed MNIST combines MNIST LeCun and Cortes (2010) and FashionMNIST Xiao et al. (2017), dividing samples into two client groups (one per dataset) with 10 clients each. Training samples were evenly distributed within each group, and the default test sets served as evaluation benchmarks. An equal number of images were generated using the global model for comparison.
CHARM is a more diverse dataset combining five domains: Cartoon faces Churchill (2019), Human faces Karras et al. (2018), Animals Xian et al. (2019), Remote sensing images Helber et al. (2019), and Marine vessels Gundogdu et al. (2016), using preprocessed square images from Meta-Album for AwA2 and MARVEL. Images were resized to $32\times 32$, and each domain was represented by 20 clients, with 20,000 images for training and 5,000 for evaluation. As with Mixed MNIST, the global model generated evaluation samples.
For Mixed MNIST, encoders and decoders used Multi-Layer Perceptrons (MLPs). On CHARM, encoders $q(z_{1}|x)$ and decoders $p(x|z_{1})$ were convolutional, while $q(z_{2}|z_{1})$ and $p(z_{1}|z_{2})$ used MLPs. Client participation followed a Bernoulli distribution: $B(0.5)$ for Mixed MNIST and $B(0.25)$ for CHARM. Hyperparameters included learning rates of $1\times 10^{-3}$ (Mixed MNIST) and $1\times 10^{-4}$ (CHARM), with 70 and 500 training rounds, respectively. Clients performed 5 local epochs per round with a batch size of 32. Centralized settings used 70 epochs for Mixed MNIST and 250 for CHARM.
Evaluation metrics included Fréchet Inception Distance Heusel et al. (2017) and Inception Score Salimans et al. (2016) for generation quality, and the negative log-likelihood (NLL) of the ELBO for reconstruction performance. IS was computed using an ImageNet-pretrained Inception model Szegedy et al. (2016).